OpenStudy (samara8954):

A CD usually sells for $15.00. If the CD is 20% off, and sales tax is 7%, what is the total price of the CD, including tax? $12.96 $12.04 $12.84 $13.64

OpenStudy (samara8954):

@MARC_

OpenStudy (anonymous):

Hmm,what do u think? @Samara8954

OpenStudy (samara8954):

I don't know?

OpenStudy (samara8954):

what do u thing @MARC_

OpenStudy (samara8954):

think

OpenStudy (anonymous):

\[(15+(15\times\frac{ 7 }{ 100 }))=\frac{ 321 }{ 20 }\] \[\frac{ 321 }{ 20 }\times \frac{ 100-20 }{ 100 }=?\] solve it @Samara8954

OpenStudy (texaschic101):

15 - 0.20(15) = the cost of the CD after the discount. So what did you get for this part of the problem ?

OpenStudy (texaschic101):

Now you would take the answer that you got and multiply it by .07....that is the tax that will be added.

OpenStudy (texaschic101):

Then you would take the answer you got for the first part of the problem, then add that to the answer you got for the second part of the problem and that will be the total.

OpenStudy (anonymous):

The CD is $15.00 Less 20% = $15.00/100 x 20 = $3 i.e the CD is $12.00 sales tax@7% = $12 x 7/100 = 84/100 = $0.84 $12 - $0.;84 = $11.16